date Find the equation of the parabola with focus at (-1,-2) and directrix x-2y +3=0 Given .
Answer #1

Question i Given focus (hik) = (-1,-2) Directrix :- = matb 2y = x+3; y = 2 + 2 / 2 So, m=1, To find Equation of parabola ; in